OpenStudy (anonymous):

cot(-8pi) find the exact value

OpenStudy (anonymous):

Do you know what the cosine of \(8\pi\) is?

OpenStudy (anonymous):

Nope! Honestly, I have no clue what's happening in class.

OpenStudy (anonymous):

That's a problem D:

OpenStudy (anonymous):

The cosine of 8pi is 1! :D that is what my calculator told me lol

OpenStudy (anonymous):

Your calculator is quite right...though, if you're allowed to use a calculator, then what's the problem? Why not make it evaluate the cotangent for you? ^.^

OpenStudy (anonymous):

Because it gives me a decimal and the answer is looking for a fraction. My calculator won't turn it into a fraction because it has a radical in the fraction.

OpenStudy (anonymous):

Let me tell you a little secret... \[\large \cot(x) = \frac1{\tan(x)}\] So if you want the cotangent of something, just divide 1 by its tangent. Go on, give Mr. Calculator another go :)

OpenStudy (anonymous):

2.5E12. It won't make fractions :(

OpenStudy (anonymous):

ooohh... Okay, what's the tangent, then?

OpenStudy (anonymous):

When I put in \[\cos (-8\pi)\div \sin(-8\pi)\]

OpenStudy (anonymous):

I get the same answer

OpenStudy (anonymous):

tangent of \(-8\pi\) what is it?

OpenStudy (anonymous):

2.5E12

OpenStudy (anonymous):

tangent. tangent. tangent. :D

OpenStudy (anonymous):

Oh mybad! Haha. 4E-13

OpenStudy (anonymous):

No... what kind of calculator are you using? D:

OpenStudy (anonymous):

TI-84

OpenStudy (anonymous):

jeez XD Why don't you just google tan(-8pi) :D

OpenStudy (anonymous):

I've tried :( it doesn't have anything for it! That's why I'm here lol. Would you happen to know tan(115pi)

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

...yes. But I'm not telling you :P You'll need to figure it out for yourself... but I will help you :) Trust me, google tan(-8pi) the answer is simpler than any fraction you could think of...

OpenStudy (anonymous):

Oh my gosh...It's 0...so the answer is undefined because it would be 1/0!

OpenStudy (anonymous):

Exactly :P See? You CAN do it ^.^

OpenStudy (anonymous):

You're smarter than your bloody TI-84 LOL

OpenStudy (anonymous):

I have no clue why it's giving my crazy answers.

OpenStudy (anonymous):

Don't trust it for now. Want to know another secret that involves your precious tan(115pi) ? ^.^

OpenStudy (anonymous):

Yes! I do enjoy a good secret!

OpenStudy (anonymous):

Well then, now that we know that Google is a bit more reliable than Mr. TI-84, why don't you ask google the following: tan(0pi) tan(1pi) tan(2pi) tan(3pi) I'll wait ^.^

OpenStudy (anonymous):

They're all 0! So tan(xpi) will always be zero! Awesome to know! Why is that? If it is a lot to explain then don't worry about it; I'll just google it.

OpenStudy (anonymous):

Correction: tan(xpi) will be zero AS LONG as x is an integer. Don't forget details, silly :P

OpenStudy (anonymous):

But yes, you have made a good observation, and the explanation isn't long at all ^.^

OpenStudy (anonymous):

And here's why: \[\Large \sin(n\pi)= 0\] for as long as n is an integer and... \[\Large \cos(n\pi) = (-1)^n\] That said, we have \[\Large \tan(n\pi) = \frac{\sin(n\pi)}{\cos(n\pi)}= \frac0{(-1)^n}=0\]

OpenStudy (anonymous):

I'm glad you didn't ask Google what tan(115pi) is, because somehow, at that point, Google becomes just as delusioned as TI-84 lawlz ^.^

OpenStudy (anonymous):

Yeah! I typed it into google and it gave me something way off. Thanks a ton! You're a life savior!